# MemInsight: Autonomous Memory Augmentation for LLM Agents

### v1

@@ Abstract
Large language model (LLM) agents have evolved to intelligently process information, make decisions, and interact with users or tools. A key capability is the integration of long-term memory capabilities, enabling these agents to draw upon historical interactions and knowledge. However, the growing memory size and need for semantic structuring pose significant challenges. In this work, we propose an autonomous memory augmentation approach, MemInsight, to enhance semantic data representation and retrieval mechanisms. By leveraging autonomous augmentation to historical interactions, LLM agents are shown to deliver more accurate and contextualized responses. We empirically validate the efficacy of our proposed approach in three task scenarios; conversational recommendation, question answering and event summarization. On the LLM-REDIAL dataset, MemInsight boosts persuasiveness of recommendations by up to 14%. Moreover, it outperforms a RAG baseline by 34% in recall for LoCoMo retrieval. Our empirical results show the potential of MemInsight to enhance the contextual performance of LLM agents across multiple tasks.
@@ 1 Introduction
LLM agents have emerged as an advanced framework to extend the capabilities of LLMs to improve reasoning Yao et al. (2023); Wang et al. (2024c), adaptability Wang et al. (2024d), and self-evolution Zhao et al. (2024a); Wang et al. (2024e); Tang et al. (2025). A key component of these agents is their memory module, which retains past interactions to allow more coherent, consistent, and personalized responses across various tasks. The memory of the LLM agent is designed to emulate human cognitive processes by simulating how knowledge is accumulated and historical experiences are leveraged to facilitate complex reasoning and the retrieval of relevant information to inform actions Zhang et al. (2024). However, the advantages of an LLM agent’s memory also introduce notable challenges Wang et al. (2024b). (1) As data accumulates over time, retrieving relevant information becomes increasingly challenging, especially during extended interactions or complex tasks. (2) Processing large historical data, which can grow rapidly, as interactions accumulate, requires effective memory management strategies. (3) Storing data in its raw format can hinder efficient retrieval of pertinent knowledge, as distinguishing between relevant and irrelevant details becomes more challenging, potentially leading to noisy or imprecise information that compromises the agent’s performance. Furthermore, (4) the integration of knowledge across tasks is constrained, limiting the agent’s ability to effectively utilize data from diverse contexts. Consequently, effective knowledge representation and structuring of LLM agent memory are essential to accumulate relevant information and enhance understanding of past events. Improved memory management enables better retrieval and contextual awareness, making this a critical and evolving area of research.
@@ 2 Related Work
Recent research in LLM agents memory focuses on developing methods for effectively storing previous interactions and feedback Packer et al. (2024). Contemporary approaches emphasize memory structures that enhance the adaptability of agents and improve their ability to generalize to previously unseen environments Zhao et al. (2024a); Zhang et al. (2024); Zhu et al. (2023). Common memory forms include summaries and abstract high-level information from raw observations to capture key points and reduce information redundancy Maharana et al. (2024). Other approaches include structuring memory as summaries, temporal events, or reasoning chains Zhao et al. (2024a); Zhang et al. (2024); Zhu et al. (2023); Maharana et al. (2024); Anokhin et al. (2024); Liu et al. (2023a). In addition, there are studies that enrich raw conversations with semantic representations like sequence of events and historical event summaries Zhong et al. (2023); Maharana et al. (2024) or extract reusable workflows from canonical examples and integrate them into memory to assist test-time inference Wang et al. (2024f). However, all aforementioned studies rely on either unstructured memory or human-designed attributes for memory representation, while MemInsight leverages the AI agent’s autonomy to discover the ideal attributes for structured representation.
@@
Existing works have leveraged memory retrieval techniques for efficiency when tackling vast amounts of historical context Hu et al. (2023a); Zhao et al. (2024b); Tack et al. (2024); Ge et al. (2025). Common approaches for memory retrieval include generative retrieval models, which encode memory as dense vectors and retrieve the top-$k$ relevant documents based on similarity search techniques Zhong et al. (2023); Penha et al. (2024). Various similarity metrics, such as cosine similarity Packer et al. (2024), are employed, alongside advanced techniques like dual-tower dense retrieval models, which encode each memory history into embeddings indexed by FAISS Johnson et al. (2017) to enhance retrieval efficiency Zhong et al. (2023). Additionally, methods such as Locality-Sensitive Hashing (LSH) are utilized to retrieve tuples containing related entries in memory Hu et al. (2023b).
@@ 3 Autonomous Memory Augmentation
Figure 1: Main modules of MemInsight including, Attribute Mining, Memory Retrieval, and Annotation, triggered by different memory processes: Augment, Write, and Retrieve. In addition to the test-time inference evaluation downstream tasks, memory augmentation and adopted memory retrieval methods.
Our proposed MemInsight model is designed to enhance memory representation through a structured augmentation process that optimizes memory retrieval. Figure 1 presents an overview of the model, highlighting its main modules: attribute mining, annotation, and memory retriever.
@@ 3.1 Attribute Mining and Annotation
To ensure the effectiveness of these attributes in future interactions, they must be meaningful, accurate, and Attribute mining in our MemInsight model leverages a backbone LLM to autonomously identify and define key attributes that encapsulate semantic knowledge from user interactions. This entails selecting attributes most relevant to the task under consideration and employs them to annotate historical conversations. Effective attributes must be meaningful, accurate, and contextually relevant to enhance future interactions. To achieve this, the augmentation process follows a structured approach, defining the perspective from which the attributes are derived, determining the appropriate level of augmentation granularity, and establishing a coherent sequence for annotation. The resulting attributes and values are then used to enrich memory, ensuring a well-organized and informative representation of past interactions.
@@ 3.1.1 Attribute Perspective
Attribute generation is guided by two primary orientations: entity-centric and conversation-centric. Entity-centric emphasizes a specific item stored in memory such as movies or books. Attributes generated for entity-centric augmentations should capture the main characteristics and features of this entity. For example, attributes for a movie entity might include the director, actors, and year of release, while attributes for a book entity would encompass the author, publisher, and number of pages. On the other hand, conversation-centric augmentations focus on annotating and characterizing the entire user interaction from the user’s perspective. This approach ensures that the extracted attributes align with the user’s intent, preferences, sentiment, emotions, motivations, and choices, thereby improving personalized responses and memory retrieval. An illustrative example is provided in Figure 4.
@@ 3.1.2 Attribute Granularity
While entity-centric augmentations focus on specific entities in memory, conversation-centric augmentations introduce an additional factor: attribute granularity, which determines the level of details captured in the augmentation process. The augmentation attributes can be analyzed at varying levels of abstraction, either at the level of individual turns within a user conversation (turn-level), or across the entire dialogue session (session-level), each offering distinct insights into the conversational context. At the turn level, each dialogue turn is independently augmented, focusing on the specific content of individual turns to generate more nuanced and contextual attributes. In contrast, session-level annotation considers the entire dialogue, generating generalized attributes that capture the broader conversational context. Due to its broader granularity, session-level augmentation emphasizes high-level attributes and conversational structures rather than the detailed features of individual turns. An example of both levels is illustrated in Figure 2 for a sample dialogue turns. As shown, turn-level annotations offer finer-grained details, while session-level annotations provide a broader overview of the dialogue.
@@ 3.1.3 Annotation and Attribute Prioritization
Subsequently, the generated attributes and their corresponding values are used to annotate the agent’s memory. Annotation is done by aggregating attributes and values in the relevant memory in the form:
$\{m_{i}:\langle a_{1},v_{1}\rangle,...,\langle a_{n},v_{n}\rangle\}$ (1)
where $m_{i}$ stands for relevant memory and $a_{i},v_{i}$ denote attributes and values respectively. The relevant memory may correspond to turn or session level. Attributes are typically aggregated using the Attribute Prioritization method, which can be classified into Basic and Priority. In Basic Augmentation, attributes are aggregated without a predefined order, resulting in an arbitrary sequence $i_{1},..,i_{n}$. In contrast, Priority Augmentation sorts attribute-value pairs according to their relevance to the memory being augmented. This prioritization follows a structured order in which attribute $i_{1}$ holds the highest significance, ensuring that more relevant attributes are processed first.
@@ 3.2 Memory Retrieval
MemInsight augmentations are employed to enrich or retrieve relevant memory. For comprehensive retrieval, memory is retrieved along with all associated augmentations to generate a more context-aware response. Additionally, MemInsight can refine the retrieval process. Initially, the current context is augmented to identify task-specific and interaction-related attributes, which then guide the retrieval of the pertinent memory. Two primary retrieval methods are proposed: (1) Attribute-based Retrieval, leverages the current context to generate attributes tailored to the specific task at hand. These attributes serve as criteria for selecting and retrieving relevant memory that shares similar attributes in their augmentations. The retrieved memories, which align with the required attributes, are subsequently integrated into the current context to enrich the ongoing interaction. (2) Embedding-based Retrieval, utilizes memory augmentations to create a unique embedding representation for each memory instance, derived from its aggregated annotations. Simultaneously, the augmentations of the current context are embedded to form a query vector, which is then used in a similarity-based search to retrieve the top-$k$ most relevant memories. Finally, all retrieved memory are incorporated into the current context to enhance the relevance and coherence of the ongoing interaction. A detailed description of this method can be found in Appendix C.
@@ 4.1 Datasets
We conduct a series of experiments on the datasets: LLM-REDIAL Liang et al. (2024) and LoCoMo Maharana et al. (2024). LLM-REDIAL is a dataset for evaluating movie Conversational Recommendation, containing approximately 10K dialogues covering 11K movies in memory. While LoCoMo is a dataset for evaluating Question Answering and Event Summarization, consisting of 30 long-term dialogues across up to 10 sessions between two speakers. LoCoMo includes five question categories: Single-hop, Multi-hop, Temporal reasoning, Open-domain knowledge, and Adversarial questions. Each question has a reference label that specifies the relevant dialogue turn in memory required to generate the answer. Additionally, LoCoMo provides event labels for each speaker in a session, which we use as ground truth for Event Summarization evaluation.
@@ 4.2 Experimental Setup
To evaluate our model, we begin by augmenting the datasets using a backbone LLM with zero-shot prompting to identify relevant attributes and their corresponding values. For augmentation generation and evaluation across various tasks, we utilize the following models for attribute generation: Claude Sonnet,111claude-3-sonnet-20240229-v1 Llama222 llama3-70b-instruct-v1 and Mistral.333mistral-7b-instruct-v0 For the Event Summarization task, we also use the Claude-3-Haiku model.444claude-3-Haiku-20240307-v1 For embedding-based retrieval tasks, we employ the Titan Text Embedding model 555titan-embed-text-v2:0 to generate embeddings. The augmented memory is then embedded and indexed using FAISS Johnson et al. (2017) for vector indexing and search. To ensure consistency across all experiments, we use the same base model for the primary tasks: recommendation, answer generation, and summarization, while evaluating different models for augmentation. Claude Sonnet serves as the backbone LLM in baselines for all tasks.
@@ 4.3 Evaluation Metrics
The evaluation metrics used for assessing different tasks using MemInsight include, traditional metrics like F1-score metric for answer prediction and recall for accuracy in Question Answering. Recall@K and NDCG@K for Conversational Recommendation, along with LLM-based metrics for genre matching.
We also evaluate using subjective metrics including Persuasiveness, used in Liang et al. (2024), to assess how persuasive the recommendations are relative to the ground truth. Additionally, we introduce a Relatedness metric, where we prompt an LLM to measure how comparable the recommendation attributes are to the ground truth, categorizing them as not comparable, comparable, or highly comparable. Finally, we assess Event Summarization using an LLM-based metric, G-Eval Liu et al. (2023b), a summarization evaluation metric that measures the relevance, consistency, and coherence of generated summaries as opposed to reference labels. These metrics provide a comprehensive framework for evaluating both retrieval effectiveness and response quality.
@@ 5 Experiments
Baseline (Claud-3-Sonnet) 15.0 10.0 3.3 26.0 45.3 26.1
@@
MemInsight (Claude-3-SonnetPriority) 15.8 15.8 6.7 19.1 75.3 30.1
@@ 5.1 Questioning Answering
Questioning Answering task experiments are conducted to evaluate the effectiveness of MemInsight in answer generation. We assess overall accuracy to measure the system’s ability to retrieve and incorporate relevant information from augmentations. The base model, which incorporates all historical dialogues without augmentation, serves as the baseline. We additionally consider Dense Passage Retrieval (DPR) RAG model  Karpukhin et al. (2020) as a comparative baseline due to its speed and scalability.
@@
In this task, memory is constructed from historical conversational dialogues, which requires the generation of conversation-centric attributes for augmentation. Given that the ground truth labels consist of dialogue turns relevant to the question, the dialogues are annotated at the turn level. A backbone LLM is prompted to generate augmentation attributes for both conversation-centric and turn-level annotations.
@@
To answer a given question, the relevant dialogue turn must be retrieved from historical dialogues. In order to retrieve the relevant dialogue turn, the question is first augmented to identify relevant attributes and a memory retrieval method is applied. We evaluate different MemInsight memory retrieval methods to demonstrate the efficacy of our model. We employ attribute-based retrieval by selecting dialogue turns augmented with attributes that exactly match the question’s attributes. Additionally, we evaluate the embedding-based retrieval, where the augmentations are embedded and indexed for retrieval. Hence, the question and attributes are transformed into an embedded query, which is used to perform a vector similarity search to retrieve the top-$k$ most similar dialogue turns. Once the relevant memory is retrieved, it is integrated into the current context to generate the final answer.
@@
We initiate our evaluation by assessing attribute-based memory retrieval using the Claude-3-Sonnet model. Table 1 presents the overall F1 score, measuring the accuracy of the generated answers. As shown in the table, attribute-based retrieval outperforms the baseline model by 3% in overall accuracy, with notable improvements in single-hop, temporal reasoning, and adversarial questions, which require advanced contextual understanding and reasoning. These results indicate that the augmented history enriched the context, leading to better reasoning and a significant increase in the F1 score for answer generation. Additionally, we perform a detailed analysis of embedding-based retrieval, where we consider evaluating basic and priority augmentation using the Claude-3-Sonnet model.
Table 1 demonstrates that the priority augmentation consistently outperforms the basic model across all questions. This finding suggests that the priority relevance of augmentations enhances context representation for conversational data. Subsequently, we evaluate the priority augmentations using Llama, and Mistral models for Embedding-based retrieval. As shown in the table, the Embedding-based retrieval outperforms the RAG baseline across all question categories, except for adversarial questions, yet the overall accuracy of MemInsight remains superior. Additionally, MemInsight demonstrates a significant improvement in performance on multi-hop questions, which require reasoning over multiple pieces of supporting evidence. This suggests that the generated augmentations provided a more robust understanding and a broader perspective of the historical dialogues. RECALL metrics in Table 2 revealed a more significant boost, with priority augmentations increasing accuracy across all categories and yielding a 35% overall improvement.

### v2

@@ Abstract
Large language model (LLM) agents have evolved to intelligently process information, make decisions, and interact with users or tools. A key capability is the integration of long-term memory capabilities, enabling these agents to draw upon historical interactions and knowledge. However, the growing memory size and need for semantic structuring pose significant challenges. In this work, we propose an autonomous memory augmentation approach, MemInsight, to enhance semantic data representation and retrieval mechanisms. By leveraging autonomous augmentation to historical interactions, LLM agents are shown to deliver more accurate and contextualized responses. We empirically validate the efficacy of our proposed approach in three task scenarios; conversational recommendation, question answering and event summarization. On the LLM-REDIAL dataset, MemInsight boosts persuasiveness of recommendations by up to 14%. Moreover, it outperforms a RAG baseline by 34% in recall for LoCoMo retrieval. Our empirical results show the potential of MemInsight to enhance the contextual performance of LLM agents across multiple tasks
@@ 1 Introduction
LLM agents have emerged as an advanced framework to extend the capabilities of LLMs to improve reasoning Yao et al. (2023); Wang et al. (2024c), adaptability Wang et al. (2024d), and self-evolution Zhao et al. (2024a); Wang et al. (2024e); Tang et al. (2025). A key component of these agents is their memory module, which retains past interactions to allow more coherent, consistent, and personalized responses across various tasks. The memory of the LLM agent is designed to emulate human cognitive processes by simulating how knowledge is accumulated and historical experiences are leveraged to facilitate complex reasoning and the retrieval of relevant information to inform actions Zhang et al. (2024). However, the advantages of an LLM agent’s memory also introduce notable challenges Wang et al. (2024b). As interactions accumulate over time, retrieving relevant information becomes increasingly difficult, especially in long-term or complex tasks. Raw historical data grows rapidly and, without effective memory management, can become noisy and imprecise, hindering retrieval and degrading agent performance. Moreover, unstructured memory limits the agent’s ability to integrate knowledge across tasks and contexts. Therefore, structured knowledge representation is essential for efficient retrieval, enhancing contextual understanding, and supporting scalable long-term memory in LLM agents. Improved memory management enables better retrieval and contextual awareness, making this a critical and evolving area of research.
@@ 2 Related Work
Recent research in LLM agents memory focuses on storing and retrieving prior interactions to improve adaptability and generalization Packer et al. (2024); Zhao et al. (2024a); Zhang et al. (2024); Zhu et al. (2023). Common approaches structure memory as summaries, temporal events, or reasoning chains to reduce redundancy and highlight key information Maharana et al. (2024); Anokhin et al. (2024); Liu et al. (2023a). Some methods enrich raw dialogues with semantic annotations, such as event sequences Zhong et al. (2023); Maharana et al. (2024) or reusable workflows Wang et al. (2024f). Recent models like A-MemXu et al. (2025) that uses manually defined task-specific notes to structure an agent’s memory, while Mem0Chhikara et al. (2025) offers a scalable, real-time memory pipeline for production use. However, most existing methods rely on unstructured memory or manually defined schemas. In contrast, MemInsight autonomously discovers semantically meaningful attributes, enabling structured memory representation without human-crafted definitions.
@@
[…]
@@ 3 Autonomous Memory Augmentation
Figure 1: MemInsight framework comprising three core modules: Attribute Mining (including perspective and granularity), Annotation (with attribute prioritization), and Memory Retrieval (including refined and comprehensive retrieval). These components are triggered by various downstream tasks such as Question Answering, Event Summarization, and Conversational Recommendation.
Our proposed model, MemInsight, encapsulates the agent’s memory $M$, offering a unified framework for augmenting and retrieving user–agent interactions represented as memory instances $m$. As new interactions occur, they are autonomously augmented and incorporated into memory, forming an enriched set $M=\{m_{1{<augmented>}},\dots,m_{n{<augmented>}}\}$. As shown in Figure 1, MemInsight comprises three core modules: attribute mining, annotation, and memory retrieval.
@@ 3.1 Attribute Mining and Annotation
Attribute mining extracts structured and semantically meaningful attributes from input dialogues for memory augmentation. The process follows a principled framework guided by three key dimensions:
(1) Perspective, from which attributes are derived (e.g., entity- or conversation-centric annotations)
(2) Granularity, indicating the level of annotation detail (e.g., turn-level or session-level)
(3) Annotation, which ensures that extracted attributes are appropriately aligned with the corresponding memory instance. A backbone LLM is leveraged to autonomously identify and generate relevant attributes.
@@ 3.1.1 Attribute Perspective
An attribute perspective entails two main orientations: entity-centric and conversation-centric. The entity-centric focuses on annotating specific items referenced in memory, such as books or movies, using attributes that capture their key properties (e.g., director, author, release year). In contrast, the conversation-centric perspective captures attributes that reflect the overall user interaction with respect to users’ intent, preferences, sentiment, emotions, motivations, and choices, thereby improving response generation and memory retrieval. An illustrative example is provided in Figure 4.
@@ 3.1.2 Attribute Granularity
Conversation-centric augmentations introduce the notion of attribute granularity, which defines the level of details captured in the augmentation process. The augmentation attributes can be analyzed at varying levels of abstraction, either at the level of individual turns within a user conversation (turn-level), or across the entire dialogue session (session-level), each offering distinct insights into the conversational context. Turn-level focuses on the specific content of individual turns to generate more nuanced and contextual attributes, while session-level augmentation captures broader patterns and user intent across the interaction. Figure 2 illustrates this distinction, showing how both levels offer complementary perspectives on a sample dialogue.
@@ 3.1.3 Annotation and Attribute Prioritization
Subsequently, the generated attributes and their corresponding values are used to annotate the agent’s memory. Annotation is done by aggregating attributes and values in the relevant memory.
Given an interaction $i$, the module applies an LLM-based extraction function $\mathcal{F}_{\text{LLM}}$ to produce a set of attribute–value pairs:
$A=\mathcal{F}_{\text{LLM}}(D)=\{(a_{j},v_{j})\}_{j=1}^{k}$
where: $a_{j}\in\mathcal{A}$ represents the attribute (e.g., emotion, entity, intent) and $v_{j}\in\mathcal{V}$ the value of this attribute These attributes are then used to annotate the corresponding memory instance $m_{i}$, resulting in an augmented memory $M_{a}$: $M_{a}=\{(A_{1},\tilde{m}_{1}),(A_{2},\tilde{m}_{2}),\dots,(A_{i},\tilde{m}_{i})\},$ Attributes are typically aggregated using the Attribute Prioritization method, which can be classified into Basic and Priority. In Basic Augmentation, attributes are aggregated without a predefined order, resulting in an arbitrary sequence. In contrast, Priority Augmentation sorts attribute-value pairs according to their relevance to the memory being augmented. This prioritization follows a structured order in which attribute $(A_{1},\tilde{m_{1}})$ holds the highest significance, ensuring that more relevant attributes are processed first.
@@ 3.2 Memory Retrieval
MemInsight augmentations are employed to both enrich memory representations and support the retrieval of contextually relevant memory. These augmentations are utilized in one of two ways.
(1) Comprehensive retrieval, retrieves all related memory instances along with their associated augmentations to support context-aware inference.
(2) Refined retrieval, where the current context is augmented to extract task-specific attributes, which then guide the retrieval process through one of the following methods:
a- Attribute-based Retrieval: which uses the current attributes as filters to select memory instances with matching or related augmentations only. Given a query session $Q$ with attributes $A_{Q}$, retrieve relevant memories:
$\mathcal{R}_{\text{attr}}(A_{Q},\mathbb{M})=\text{Top-}k\left\{(A_{k},M_{k})\mid\text{match}(A_{Q},A_{k})\right\}$
[…]
$\mathcal{R}_{\text{embed}}(A_{Q},\mathbb{M})=\text{Top-}k\left\{(A_{k},M_{k})\mid\text{sim}(A_{Q},A_{k})\right\}$
Finally, the retrieved memories are then integrated into the current context to inform the ongoing interaction. Further implementation details of embedding-based retrieval are provided in Appendix C.
@@ 4.1 Datasets
We evaluate MemInsight on two benchmarks: LLM-REDIAL Liang et al. (2024) and LoCoMo Maharana et al. (2024). LLM-REDIAL is a dataset for conversational movie recommendation, comprising  10K dialogues and 11K movie mentions. LoCoMo is a dataset for evaluating Question Answering and Event Summarization, with 30 multi-session dialogues between two speakers. It features five question types: Single-hop, Multi-hop, Temporal reasoning, Open-domain, and Adversarial, each annotated with the relevant dialogue turn required for answering. LoCoMo also provides event labels for each speaker in a session, which serve as ground truth for evaluating event summarization.
@@ 4.2 Experimental Setup
To evaluate our model, we begin by augmenting the datasets using zero-shot prompting to extract relevant attributes and their corresponding values. For attribute generation across tasks, we employ Claude Sonnet111claude-3-sonnet-20240229-v1, LLaMA 3222llama3-70b-instruct-v1, and Mistral333mistral-7b-instruct-v0. For the Event Summarization task, we additionally utilize Claude 3 Haiku444claude-3-haiku-20240307-v1. In embedding-based retrieval, we use the Titan Text Embedding model555titan-embed-text-v2:0 to generate embeddings of augmented memory, which are indexed and searched using FAISS Johnson et al. (2017). To ensure consistency across all experiments, we use the same base model for the primary tasks: recommendation, answer generation, and summarization, while varying the models used for memory augmentation. Claude Sonnet serves as the backbone LLM in all baseline evaluations.
@@ 4.3 Evaluation Metrics
We evaluate MemInsight using a combination of standard and LLM-based metrics. For Question Answering, we report F1-score for answer prediction and recall for accuracy; for Conversational Recommendation, we use Recall@K, NDCG@K, along with LLM-based metrics for genre matching.
We further incorporate subjective metrics, including Persuasiveness  Liang et al. (2024), which measures how persuasive a recommendation aligns with the ground truth. Additionally, we introduce a Relatedness metric where we prompt an LLM to measure how comparable are recommendation attributes to the ground truth, categorizing them as not comparable, comparable, or highly comparable. For Event Summarization, we adopt G-Eval Liu et al. (2023b), an LLM-based metric that evaluates the relevance, consistency, and coherence of generated summaries against reference labels. Together, these metrics provide a comprehensive framework for evaluating both retrieval effectiveness and response quality.
@@ 5 Experiments
Baseline (Claude-3-Sonnet) 15.0 10.0 3.3 26.0 45.3 26.1
LoCoMo (Mistral v1) 10.2 12.8 16.1 19.5 17.0 13.9
ReadAgent (GPT-4o) 9.1 12.6 5.3 9.6 9.81 8.5
MemoryBank (GPT-4o) 5.0 9.6 5.5 6.6 7.3 6.2
@@
MemInsight (Claude-3-SonnetPriority) 15.8 15.8 6.7 19.7 75.3 30.1
@@ 5.1 Questioning Answering
Question Answering experiments are conducted to evaluate the effectiveness of MemInsight in answer generation. We evaluate the overall accuracy to measure the system’s ability to retrieve and integrate relevant information using memory augmentations. The base model, which incorporates all historical dialogues without any augmentation, serves as a baseline. Additionally, we report results on the LoCoMo benchmark using the same backbone model (Mistral v1) to ensure a fair evaluation. We also compare with stronger GPT-based baselines, including MemoryBankZhong et al. (2023) and ReadAGent Lee et al. (2024), which utilizes external memory modules to support long-term reasoning. We also consider Dense Passage Retrieval (DPR) Karpukhin et al. (2020) as a representative baseline of RAG due to its scalability and retrieval efficiency.
@@
In this task, memory is constructed from historical conversational dialogues, which requires the generation of conversation-centric attributes for augmentation. Given that the ground-truth labels consist of dialogue turns relevant to the question, the dialogues are annotated at the turn level. An LLM backbone is prompted to generate augmentation attributes for both conversation-centric and turn-level annotations.
@@
[…]
@@
As shown in Table 1, MemInsight achieves significantly higher overall accuracy on the question answering task compared to all baselines, using both attribute-based and embedding-based memory retrieval. In the attribute-based setting, MemInsight with Claude-3-Sonnet demonstrates notable gains in single-hop, temporal, and adversarial questions, which require more complex contextual reasoning. These results highlight the effectiveness of memory augmentation in enriching context and enhancing answer quality. MemInsight further outperforms all other benchmark models across most question types, with the exception of multi-hop and temporal questions in LoCoMo, where evaluation is based on a partial-match F1 metric Maharana et al. (2024).
For embedding-based retrieval, we evaluate MemInsight using both basic and priority augmentation, alongside the DPR baseline. MemInsight consistently outperforms all baselines, except in temporal and adversarial questions, where DPR achieves slightly higher accuracy. Nevertheless, MemInsight maintains the highest overall accuracy. Priority augmentation also consistently outperforms basic augmentation across nearly all question types, validating its effectiveness in improving contextual relevance. Notably, MemInsight demonstrates substantial gains on multi-hop questions, which require reasoning over multiple pieces of supporting evidence, highlighting its ability to integrate dispersed information from historical dialogue. As shown in Table 2, recall metrics further support this trend, with priority augmentation yielding a 35% overall improvement and consistent gains across all categories.